# A Linear Quantum Coupler for Clean Bosonic Control

### v1

@@ Abstract
Quantum computing with superconducting circuits relies on high-fidelity driven nonlinear processes. An ideal quantum nonlinearity would selectively activate desired coherent processes at high strength, without activating parasitic mixing products or introducing additional decoherence. The wide bandwidth of the Josephson nonlinearity makes this difficult, with undesired drive-induced transitions and decoherence limiting qubit readout, gates, couplers and amplifiers. Significant strides have been recently made into building better ‘quantum mixers’, with promise being shown by Kerr-free three-wave mixers that suppress driven frequency shifts, and balanced quantum mixers that explicitly forbid a significant fraction of parasitic processes. We propose a novel mixer that combines both these strengths, with engineered selection rules that make it essentially linear (not just Kerr-free) when idle, and activate clean parametric processes even when driven at high strength. Further, its ideal Hamiltonian is simple to analyze analytically, and we show that this ideal behavior is first-order insensitive to dominant experimental imperfections. We expect this mixer to allow significant advances in high-Q control, readout, and amplification.
@@ I Introduction
Driven superconducting circuits provide a platform for studying complex nonlinear quantum optics and manipulating quantum information in the microwave domain. The intrinsic nonlinearity of Josephson junctions [1] has been utilized to design numerous circuit elements, ranging from quantum-limited amplifiers [2, 3, 4, 5, 6, 7, 8, 9], novel non-reciprocal devices [10, 11, 12, 13], ultra-low-noise detectors [14, 15, 16], and, most notably, quantum information processing through the field of circuit Quantum Electrodynamics [17, 18]. A significant portion of these applications are enabled by nonlinear quantum mixers that activate specific desired processes when parametrically driven with a microwave drive. Such parametric mixers have found use as couplers [19, 20, 21, 22], amplifiers [9], and even nonlinear oscillators in their own right [23, 24].
@@
Figure 1: The LINC, a quantum single-balanced mixer a, Fundamentals of mixer balancing. By taking two identical nonlinear elements and driving them $180^{\circ}$ out of phase (e.g., with a balun), we can create a single-balanced mixer. If the nonlinear element is a single junction, this circuit is the familiar differentially-driven SQUID (DDS, [21]). If the nonlinear element is instead a DC flux-biased RF-SQUID, this circuit results in the Linear Inductive Coupler (LINC). b, An illustration of the advantage of a balanced mixer. In general, a driven nonlinear element will permit numerous nonlinear processes, of which only a small subset is typically desired. Balancing the nonlinear element will suppress a significant fraction of spurious processes (light orange), but preserve the desired process (black) at high strength. Any remaining parasitic processes (red) still permitted by the symmetry must be avoided through other means, like careful selection of drive frequencies. c, The LINC enables a separation of driven and undriven mixer behavior. At the operating point where the total DC flux in the outer SQUID loop is half flux quantum, the loop junctions are biased to effectively infinite inductance, leaving only the linear shunt as the coupler’s inductor. This simultaneously nulls all nonlinearity when idle, which is beyond just being Kerr-free. When threaded with AC flux, the outer loop activates and drives a balanced three-wave mixing process.
@@ II The Linear Inductive Coupler
Let us now analyze the circuit’s behavior, both when idle and when driven. At the special flux point of $\phi_{\text{DC}}=\frac{\pi}{2}$, the nonlinearity of the circuit completely disappears when idle, resulting in a completely linear static Hamiltonian:
@@
This is beyond a simple ‘Kerr-free’ mixer – at this flux point, the circuit’s static nonlinearity at all orders completely disappears. We will show later that this property holds even in the presence of experimental imperfections like a parasitic linear inductance, which is often present in realistic mixers due to geometric constraints. The mixer hence provides a completely linear environment to any modes it mediates a coupling between in the coupler context, which is particularly useful for bosonic control. We thus name this mixer the ‘Linear INductive Coupler’ (LINC).
@@
We note that while the simultaneous balancing of the drive and the bias of this circuit is novel and crucial to its function, the circuit topology itself is fairly common in literature [20, 36, 37, 38, 39]. The LINC is a sibling to the Asymmetrically Threaded SQUID (ATS), with nominally identical circuit geometries but an exactly opposing notion of bias point. This allows the LINC to be optimized as a three-wave mixing coupler, while the ATS is utilized for Kerr-free four-wave mixing applications, like parametric two-photon dissipation [20]. The LINC also shares its topology with the gradiometric SNAIL, where the flux bias in one of the two loops is used to tune the effective Hamiltonian of a regular charge-driven SNAIL, and symmetry is not a strict constraint. One key physical feature where the LINC might differ from these circuits is in its shunting inductor, which for the SNAIL and ATS consist of an array of a few junctions. Since bosonic applications might require highly linear idle circuits, the LINC shunt could benefit from strategies like using high kinetic inductance materials such as NbN or granular Aluminum (grAl) or a meandering geometric inductor, to minimize its static nonlinearity.
@@ IV The LINC as a driven mixer
The dominant mixing process in the driven LINC is an effective 3-wave mixing process, which for small drive strengths is given by:
$\begin{split}H_{\text{driven}}^{(3)}&=\left(\dfrac{E_{J}}{2E_{L}}\right)\omega_{L}\;J_{1}\left(|\phi_{\text{AC}}|\right)\cos{\omega_{d}t}\left(c+c^{\dagger}\right)^{2}\\ &\approx\left.g_{\text{3wm}}\right|_{\frac{\pi}{2}}\;\phi_{\text{AC}}(t)\left(c+c^{\dagger}\right)^{2},\end{split}$ (9)
where $\omega_{L}=\sqrt{8E_{L}E_{C}}$ is the frequency of the LINC mode, $c$ and $c^{\dagger}$ are its ladder operators, $J_{1}(x)$ is the first-order Bessel function, and $\phi_{\text{AC}}(t)=|\phi_{\text{AC}}|\cos{\omega_{d}t}$ is the drive. For two modes Alice ($\hat{a}$, $\omega_{a}$) and Bob ($\hat{b}$, $\omega_{b}$) that are coupled to the LINC (with energy participations $p_{a}$ and $p_{b}$), driving at select frequencies can activate various desired processes, such as:
$\begin{split}\omega_{d}=\omega_{a}-\omega_{b}&\implies g_{\text{BS}}\left(a^{\dagger}b+ab^{\dagger}\right)\\ \omega_{d}=\omega_{a}+\omega_{b}&\implies g_{\text{TS}}\left(a^{\dagger}b^{\dagger}+ab\right)\\ \omega_{d}=2\omega_{a}&\implies g_{\text{Sa}}\left({a^{\dagger}}^{2}+a^{2}\right)\\ g_{\text{BS}}=g_{\text{TS}}&=2g_{\text{Sa}}\approx g_{\text{3wm}}|\phi_{\text{AC}}|\sqrt{p_{a}p_{b}},\end{split}$ (10)
@@ V Effect of asymmetry and parasitic inductance
To quantify the asymmetry of the outer junctions, we use the ratio $(E_{J1}-E_{J2})/E_{L}\coloneqq\beta_{\Delta}$, which from fabrication imperfections we expect to be $\lesssim 2\%$. The second imperfection is a finite difference in the DC flux in the two LINC sub-loops, arising from a gradient in the residual magnetic field in which the experimental package cooled down. This imperfection is in principle possible to cancel in-situ with two dedicated flux lines per coupler, such that the relative currents in the flux lines can be tuned to achieve arbitrary flux biases in the two sub-loops. However, the LINC ideally only requires a single fast flux line for its drive and DC flux, which always applies symmetric flux to both sub-loops. In this scenario, a residual DC flux difference $\phi_{\Delta}$ will change both the LINC’s driven and undriven behavior, which we will analyze below. In the fully general case, one may also have an asymmetry in the applied drive due to imperfect drive line engineering, which might also affect performance. However, this asymmetry can be nearly entirely nulled through careful microwave design [21], so we will not explore its effect here.
@@
$K_{\text{asym}}\approx E_{c}\beta_{\Delta}\sin\left(\phi_{\Delta}-(1+\sqrt{2})\beta_{\Delta}\right)$ (14)
@@
When a LINC with non-zero asymmetry is driven, the LINC may turn on processes that would have otherwise be parity-protected. As an example, we consider the third-order mixing process involving two drive photons and a single LINC photon, corresponding to a parasitic sub-harmonic displacement of the coupler. This process is always allowed in any charge-driven mixer (like the SNAIL), yet is forbidden in the ideal LINC. The strength of this process, which we represent as $H_{\text{sub}}=g_{1,2}\phi_{\text{AC}}^{2}(c+c^{\dagger})$, is given by (appendix C):
@@ VI Conclusions and Outlook
In the bosonic context, the LINC promises to break the trade-off between fast nonlinear control and the idle errors introduced by a nonlinear ancillary mode. This is particularly important for multi-photon encodings, where inherited Kerr and thermal-noise induced dephasing irreversibly spoil logical information. Even in single-photon encodings like the dual-rail bosonic qubit, the LINC should reduces static dispersive interactions between neighboring oscillators, while enabling fast and clean universal control through parametric beamsplitting. This advantage could even extend to non-Gaussian control, where the LINC could serve as an interface between the bosonic mode and its ancillary qubit, effectively shielding the nonlinearity when idling [49].
@@ VII Acknowledgements
Figure 4: Analyzing the LINC circuit a, Modes of operation of the circuit. The circuit, with an arbitrary inductive element as its center shunt, is defined by the three variables $\left\{\theta_{1},\theta_{2},\theta_{s}\right\}$. These can be re-written in terms of the more operationally relevant charge-dipole ($\hat{\theta}_{c}$), symmetric flux ($\hat{\phi}_{sym}$), and anti-symmetric flux ($\hat{\phi}_{asym}$) modes. b, LINC three-wave mixing strength as a function of the operating point, comparing analytic formula (grey dashed, eq. 25) to exact Floquet results from a squeezing operation within the LINC ($g_{\text{SL}}$, black) and a beamsplitter operation between two external resonators (teal, details in appendix B). The discrepancy in the beamsplitting prediction may be due to driven changes in the effective resonator-LINC participations, which are captured in the Floquet simulation but not in the analytics. c, Comparison of the analytic formulae (grey dashed, eq. 27) to exact Floquet simulation for driven coupler Zeeman shift as a function of DC flux, for fixed drive amplitude $\phi_{AC}=0.1\pi$.
@@ Appendix A Analyzing the LINC circuit
$\begin{split}g_{mn}[\phi_{\text{DC}}]&:=\dfrac{1}{m!\>n!}\sum_{k=0}^{n}\binom{n}{k}\left.u_{m,n-k}\;\frac{\partial^{k}\theta_{zpf}^{m}}{\partial\phi_{d}^{k}}\right|_{\phi_{d}=\phi_{\text{DC}}},\\ u_{m,k}&:=\left.\dfrac{\partial^{m+k}\;U_{\text{LINC}}}{\partial\hat{\theta}^{m}_{c}\;\partial\phi_{d}^{k}}\right|_{\hat{\theta}_{c}=0}.\end{split}$ (22)
@@ Appendix B Driven operations with the LINC
$\begin{split}\hat{c}&=\tilde{c}+\sqrt{p_{a}}\tilde{a}+\sqrt{p_{b}}\tilde{b}\\ p_{a},p_{b}&=\left(\frac{g_{a}}{\Delta_{La}}\right)^{2},\left(\frac{g_{b}}{\Delta_{Lb}}\right)^{2},\end{split}$ (28)
@@
$\begin{split}H_{\text{3wm}}&=\tilde{\omega}_{a}\tilde{a}^{\dagger}\tilde{a}+\tilde{\omega}_{b}\tilde{b}^{\dagger}\tilde{b}+\tilde{\omega}_{L}\tilde{c}^{\dagger}\tilde{c}\\ &+g_{\text{3wm}}\frac{|\phi_{\text{AC}}|}{2}\left(e^{i\omega_{d}t}+e^{-i\omega_{d}t}\right)\times\\ &\left(\sqrt{p_{a}}\;\tilde{a}+\sqrt{p_{b}}\;\tilde{b}+\tilde{c}+\text{h.c.}\right)^{2}\end{split}$ (29)
@@
$\begin{split}H_{\text{LINC}}&=\omega_{a}a^{\dagger}a+\omega_{a}b^{\dagger}b+\omega_{c}c^{\dagger}c\\ &-g_{ac}(a^{\dagger}-a)(c^{\dagger}-c)-g_{bc}(b^{\dagger}-b)(c^{\dagger}-c)\\ &-2M^{2}E_{J}\sin{\phi_{\text{AC}}}\cos{\frac{\hat{\theta}_{c}}{M}},\end{split}$ (31)
@@
$\begin{split}H_{\text{SNAIL}}&=\omega_{a}a^{\dagger}a+\omega_{a}b^{\dagger}b+\omega_{c}c^{\dagger}c\\ &-g_{ac}(a^{\dagger}-a)(c^{\dagger}-c)-g_{bc}(b^{\dagger}-b)(c^{\dagger}-c)\\ &-\alpha M^{2}E_{J}\cos_{\text{nl}}{\left(\theta_{\text{eq,1}}+\phi_{\text{AC}}+\frac{\hat{\theta}}{M}\right)}\\ &-NM^{2}E_{J}\cos_{\text{nl}}{\left(\frac{1}{N}\left(\theta_{\text{eq,2}}+\phi_{\text{AC}}+\frac{\hat{\theta}}{M}\right)\right)},\end{split}$ (32)

### v2

@@ Abstract
Quantum computing with superconducting circuits relies on high-fidelity driven nonlinear processes. An ideal quantum nonlinearity would selectively activate desired coherent processes at high strength, without activating parasitic mixing products or introducing additional decoherence. The wide bandwidth of the Josephson nonlinearity makes this difficult, with undesired drive-induced transitions and decoherence limiting qubit readout, gates, couplers, and amplifiers. Significant strides have been recently made into building better ‘quantum mixers’, with promise being shown by Kerr-free three-wave mixers that suppress driven frequency shifts, and balanced quantum mixers that explicitly forbid a significant fraction of parasitic processes. We propose a novel mixer that combines both these strengths, with engineered selection rules that make it essentially linear (not just Kerr-free) when idle, and activate clean parametric processes even when driven at high strength. Further, its ideal Hamiltonian is simple to analyze analytically, and we show that this ideal behavior is first-order insensitive to dominant experimental imperfections. We expect this mixer to allow significant advances in high-Q control, readout, and amplification.
@@ I Introduction
Driven superconducting circuits provide a platform for studying complex nonlinear quantum optics and manipulating quantum information in the microwave domain. The intrinsic nonlinearity of Josephson junctions [1] has been utilized to design numerous circuit elements, ranging from quantum-limited amplifiers [2, 3, 4, 5, 6, 7, 8, 9], novel non-reciprocal devices [10, 11, 12, 13], ultra-low-noise detectors [14, 15, 16], and, most notably, quantum information processing through the field of circuit Quantum Electrodynamics [17, 18]. A significant portion of these applications are enabled by nonlinear quantum mixers that activate specific desired processes when parametrically driven with a microwave drive. Such parametric mixers have found use as couplers [19, 20, 21, 22], amplifiers [9], and even nonlinear oscillators in their own right [23, 24, 25].
@@
Figure 1: The LINC, a quantum single-balanced mixer a, Fundamentals of mixer balancing. By taking two identical nonlinear elements and driving them $180^{\circ}$ out of phase (e.g., with a balun), we can create a single-balanced mixer. If the nonlinear element is a single junction, this circuit is the familiar differentially-driven SQUID (DDS, [21]). If the nonlinear element is instead a DC flux-biased RF-SQUID, this circuit results in the Linear Inductive Coupler (LINC). b, An illustration of the advantage of a balanced mixer. In general, a driven nonlinear element will permit numerous nonlinear processes, of which only a small subset is typically desired. Balancing the nonlinear element will suppress a significant fraction of spurious processes (light orange), but preserve the desired process (black) at high strength. Any remaining parasitic processes (red) still permitted by the symmetry must be avoided through other means, like careful selection of drive frequencies. c, The LINC circuit, with a symmetric loop and capacitive pads, enables a separation of driven and undriven mixer behavior. At the operating point where the total DC flux in the outer SQUID loop is half a flux quantum, the loop junctions are biased to effectively infinite inductance, leaving only the linear shunt as the coupler’s inductor. This simultaneously nulls all nonlinearity when idle, which is beyond just being Kerr-free. When threaded with AC flux, the outer loop activates and drives a balanced three-wave mixing process.
@@ II The Linear Inductive Coupler
Let us now analyze the circuit’s behavior, both when idle and when driven. At the special flux point of $\phi_{\text{DC}}=\frac{\pi}{2}$, the nonlinearity of the circuit completely disappears when idle, resulting in an exactly linear static Hamiltonian:
@@
This is beyond a simple ‘Kerr-free’ mixer – at this flux point, the circuit’s static nonlinearity at all orders vanishes. We will show later that this property holds even in the presence of experimental imperfections like a parasitic series linear inductance, which is often present in realistic mixers due to geometric constraints. The mixer hence provides an extremely linear environment to any modes it mediates a coupling between in the coupler context, which is particularly useful for bosonic control. We thus name this mixer the ‘Linear INductive Coupler’ (LINC).
@@
We note that while the simultaneous balancing of the drive and the bias of this circuit is novel and crucial to its function, the circuit topology itself is fairly common in literature [20, 37, 38, 39, 40]. The LINC is a sibling to the Asymmetrically Threaded SQUID (ATS), with nominally identical circuit geometries but an exactly opposing notion of bias point. This allows the LINC to be optimized as a three-wave mixing coupler, while the ATS is utilized for Kerr-free four-wave mixing applications, like parametric two-photon dissipation [20]. The LINC also shares its topology with the gradiometric SNAIL [37], where the flux bias in the SQUID loops is used to tune the effective junction ratio of a regular charge-driven SNAIL, and symmetry is not a strict constraint. One key physical feature where the LINC might differ from these circuits is in its shunting inductor, which for the SNAIL and ATS typically consists of an array of a few junctions. Since bosonic applications might require highly linear idle circuits, the LINC shunt could benefit from using high kinetic inductance materials such as NbN or granular Aluminum (grAl), or a meandering geometric inductor, to minimize its static nonlinearity.
@@ IV The LINC as a driven mixer
The dominant mixing process in the driven LINC is an effective 3-wave mixing process, which for small drive strengths at $\phi_{\text{DC}}=\pi/2$ is given by:
$\begin{split}H_{\text{driven}}^{(3)}&=\left(\dfrac{E_{J}}{2E_{L}}\right)\omega_{L}\;J_{1}\left(|\phi_{\text{AC}}|\right)\cos{\omega_{d}t}\left(\hat{c}+\hat{c}^{\dagger}\right)^{2}\\ &\approx\left.g_{\text{3wm}}\right|_{\frac{\pi}{2}}\;\phi_{\text{AC}}(t)\left(\hat{c}+\hat{c}^{\dagger}\right)^{2},\end{split}$ (9)
where $\omega_{L}=\sqrt{8E_{L}E_{C}}$ is the frequency of the LINC mode at the operating point, $\hat{c}$ and $\hat{c}^{\dagger}$ are its ladder operators, $J_{1}(x)$ is the first-order Bessel function, and $\phi_{\text{AC}}(t)=|\phi_{\text{AC}}|\cos{\omega_{d}t}$ is the drive. For two modes Alice ($\hat{a}$, $\omega_{a}$) and Bob ($\hat{b}$, $\omega_{b}$) that are coupled to the LINC (with energy participations $p_{a}$ and $p_{b}$), driving at select frequencies can activate various desired processes, such as:
$\begin{split}\omega_{d}=\omega_{a}-\omega_{b}&\implies g_{\text{BS}}\left(\hat{a}^{\dagger}\hat{b}+\hat{a}\hat{b}^{\dagger}\right)\\ \omega_{d}=\omega_{a}+\omega_{b}&\implies g_{\text{TS}}\left(\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{a}\hat{b}\right)\\ \omega_{d}=2\omega_{a}&\implies g_{\text{Sa}}\left(\hat{a}^{\dagger^{2}}+\hat{a}^{2}\right)\\ g_{\text{BS}}=g_{\text{TS}}&=2g_{\text{Sa}}\approx g_{\text{3wm}}|\phi_{\text{AC}}|\sqrt{p_{a}p_{b}},\end{split}$ (10)
@@ V Effect of asymmetry and parasitic inductance
To quantify the asymmetry of the outer junctions, we use the ratio $(E_{J1}-E_{J2})/E_{L}\coloneqq\beta_{\Delta}$, which from fabrication imperfections we expect to be $\lesssim 2\%$. The second imperfection is a finite difference in the DC flux in the two LINC sub-loops, arising from a gradient in the residual magnetic field in which the experimental package cooled down. This imperfection is in principle possible to cancel in-situ with two dedicated flux lines per coupler, such that the relative currents in the flux lines can be tuned to achieve arbitrary flux biases in the two sub-loops. However, the LINC ideally only requires a single fast flux line for its drive and DC flux, which always applies symmetric flux to both sub-loops. In this scenario, a residual DC flux difference $\phi_{\Delta}$ will change both the LINC’s driven and undriven behavior, which we will analyze below. In the fully general case, one may also have an asymmetry in the applied drive due to imperfect drive line engineering, which might also affect performance. However, this asymmetry can be largely avoided by appropriately designing the flux distribution and capacitance matrix of the device, aided by numerical simulation techniques [21, 50].
@@
$\alpha_{L}^{\text{asym}}\approx E_{c}\beta_{\Delta}\sin\left(\phi_{\Delta}-(1+\sqrt{2})\beta_{\Delta}\right)$ (14)
@@
When a LINC with non-zero asymmetry is driven, the LINC may turn on processes that would have otherwise been parity-protected. As an example, we consider the third-order mixing process involving two drive photons and a single LINC photon, corresponding to a parasitic sub-harmonic displacement of the coupler. This process is always allowed in any charge-driven mixer (like the SNAIL), yet is forbidden in the ideal LINC. The strength of this process, which we represent as $H_{\text{sub}}=g_{1,2}\phi_{\text{AC}}^{2}(\hat{c}+\hat{c}^{\dagger})$, is given by (appendix C):
@@ VI Conclusions and Outlook
In the bosonic context, the LINC promises to break the trade-off between fast nonlinear control and the idle errors introduced by a nonlinear ancillary mode (similar to [51, 52, 25, 53]). This is particularly important for multi-photon encodings, where inherited Kerr and thermal-noise induced dephasing irreversibly spoil logical information. Even in single-photon encodings like the dual-rail bosonic qubit, the LINC should reduces static dispersive interactions between neighboring oscillators, while enabling fast and clean universal control through parametric beamsplitting. This advantage could even extend to non-Gaussian control, where the LINC could serve as an interface between the bosonic mode and its ancillary qubit, effectively shielding the nonlinearity when idling [54].
@@ VII Acknowledgements
Figure 4: Analyzing the LINC circuit a, Modes of operation of the circuit. The circuit, with an arbitrary inductive element as its center shunt, is defined by the three variables $\left\{\theta_{1},\theta_{2},\theta_{s}\right\}$. These can be re-written in terms of the more operationally relevant charge-dipole ($\hat{\theta}_{c}$), symmetric flux ($\hat{\phi}_{sym}$), and anti-symmetric flux ($\hat{\phi}_{asym}$) modes. b, LINC three-wave mixing strength as a function of the operating point, comparing analytic formula (grey dashed, eq. 25) to exact time-domain and Floquet results from a squeezing operation within the LINC ($g_{\text{SL}}$, black) and a beamsplitter operation between two external resonators (teal, details in appendix B) respectively. The discrepancy in the beamsplitting prediction may be due to driven changes in the effective resonator-LINC participations, which are captured in the Floquet simulation but not in the analytics. c, Comparison of the analytic formulae (grey dashed, eq. 27) to exact Floquet simulation for driven coupler Zeeman shift as a function of DC flux, for fixed drive amplitude $\phi_{AC}=0.1\pi$.
@@ Appendix A Analyzing the LINC circuit
$\begin{gathered}g_{mn}[\phi_{\text{DC}}]:=\dfrac{1}{m!\>n!}\sum_{k=0}^{n}\binom{n}{k}\left.u_{m,n-k}\;\frac{\partial^{k}\theta_{zpf}^{m}}{\partial\phi_{d}^{k}}\right|_{\phi_{d}=\phi_{\text{DC}}},\\ u_{m,k}:=\left.\dfrac{\partial^{m+k}\;U_{\text{LINC}}}{\partial\hat{\theta}^{m}_{c}\;\partial\phi_{d}^{k}}\right|_{\hat{\theta}_{c}=0}.\end{gathered}$ (22)
@@ Appendix B Driven operations with the LINC
$\begin{split}\hat{c}&=\hat{\tilde{c}}+\sqrt{p_{a}}\hat{\tilde{a}}+\sqrt{p_{b}}\hat{\tilde{b}}\\ p_{a},p_{b}&=\left(\frac{g_{a}}{\Delta_{La}}\right)^{2},\left(\frac{g_{b}}{\Delta_{Lb}}\right)^{2},\end{split}$ (28)
@@
$\begin{split}H_{\text{3wm}}&=\tilde{\omega}_{a}\hat{\tilde{a}}^{\dagger}\hat{\tilde{a}}+\tilde{\omega}_{b}\hat{\tilde{b}}^{\dagger}\hat{\tilde{b}}+\tilde{\omega}_{L}\hat{\tilde{c}}^{\dagger}\hat{\tilde{c}}\\ &+g_{\text{3wm}}\frac{|\phi_{\text{AC}}|}{2}\left(e^{i\omega_{d}t}+e^{-i\omega_{d}t}\right)\times\\ &\left(\sqrt{p_{a}}\;\hat{\tilde{a}}+\sqrt{p_{b}}\;\hat{\tilde{b}}+\hat{\tilde{c}}+\text{h.c.}\right)^{2}\end{split}$ (29)
@@
$\begin{split}H_{\text{LINC}}&=\omega_{a}\hat{a}^{\dagger}\hat{a}+\omega_{a}\hat{b}^{\dagger}\hat{b}+\omega_{c}\hat{c}^{\dagger}\hat{c}\\ &-g_{ac}(\hat{a}^{\dagger}-\hat{a})(\hat{c}^{\dagger}-\hat{c})-g_{bc}(\hat{b}^{\dagger}-\hat{b})(\hat{c}^{\dagger}-\hat{c})\\ &-2M^{2}E_{J}\sin{\phi_{\text{AC}}}\cos{\frac{\hat{\theta}_{c}}{M}},\end{split}$ (31)
@@
$\begin{split}H_{\text{SNAIL}}&=\omega_{a}\hat{a}^{\dagger}\hat{a}+\omega_{a}\hat{b}^{\dagger}\hat{b}+\omega_{c}\hat{c}^{\dagger}\hat{c}\\ &-g_{ac}(\hat{a}^{\dagger}-\hat{a})(\hat{c}^{\dagger}-\hat{c})-g_{bc}(\hat{b}^{\dagger}-\hat{b})(\hat{c}^{\dagger}-\hat{c})\\ &-\alpha M^{2}E_{J}\cos_{\text{nl}}{\left(\theta_{\text{eq,1}}+\phi_{\text{AC}}+\frac{\hat{\theta}}{M}\right)}\\ &-NM^{2}E_{J}\cos_{\text{nl}}{\left(\frac{1}{N}\left(\theta_{\text{eq,2}}+\phi_{\text{AC}}+\frac{\hat{\theta}}{M}\right)\right)},\end{split}$ (32)